# The Susceptibility of Example-Based Explainability Methods to Class Outliers

### v1

@@ 1 Introduction
Overall, our study reveals that (a) existing example-based explainability methods are not robust to class outliers, and (b) attempting to improve robustness by suppressing class outliers can lead to a loss of correctness. We propose a new evaluation framework, specifically for example-based explainers, that includes metrics for relevance, distinguishability, and correctness, and which can quantitatively assess the impact of class outliers on the effectiveness of explainers.
@@ 3.4 Explainer Correctness
Correctness quantifies the degree to which the explanations align with the underlying labeling rule. Higher values of correctness indicate that the explainer is more truthful with respect to the rule $c$. Observe that correctness is essentially the precision with which an explainer returns rule followers and breakers. We can diffentiate between correctness with respect to rule followers and breakers. Finally, we note that an important assumption is that the model $f$ is itself truthful to the rule, i.e., it has correctly learned the rule $c$, a condition we can check after training.
@@ 4.1 Relevance
To compute relevance, we employ cosine similarity, given that the dataset comprises image embeddings produced by a pre-trained model. Utilizing the graph representation, we determine the average similarity between a explanandum node (explanandum) and its corresponding neighbors (example). The first column of Table 1, shows the explainer relevance for all explainability methods. RIF demonstrates superior performance, whereas the remaining methods exhibit similar scores.
@@ 4.3 Correctness
While explaining an explanandum following the rule we expect rule-followers in the training set with the same label to have a positive influence on the prediction, and rule-breakers with the opposite label to have a negative influence. Table 1 shows that IF and DataModels perform well in terms of correctness, but are inconsistent in terms of relevance and distinguishability (domain and overlap), which is expected due to the presence of class outliers that are globally important. In contrast, RIF performs poorly in uncovering followers and breakers, because of its loss-based outlier elimination. RIF treats training data with high loss as undesired, and excludes them from explanation lists—the rationale is that such data would appear in all explanations and thus have little utility. But in this case, it is precisely the rule followers and particularly the minority of rule breakers that have high losses due to the ambiguity in the labeling rule. On the contrary, as previously discussed, the examples with very low loss are rewarded and appear in the explanations. TraceIn also fails to uncover the rule due to its low efficiency in identifying truly important examples, which is also demonstrated in [10].
@@ 5 Conclusion
Our analysis suggests that current example-based explainability techniques are susceptible to the influence of class outliers. Outright removal of outliers from explanations or datasets may not always be advisable, as they can possess valuable insights and provide explanations for similar examples encountered by models.

### v2

@@ 1 Introduction
Overall, our study reveals that (a) existing example-based explainability methods are not robust to class outliers, and (b) attempting to improve robustness by suppressing class outliers can lead to a loss of correctness. We propose a new evaluation framework, specifically for example-based explainers, including metrics for relevance, distinguishability, and correctness, and which can quantitatively assess the impact of class outliers on the effectiveness of explainers.
@@ 3.4 Explainer Correctness
Correctness quantifies the degree to which the explanations align with the underlying labeling rule. Higher correctness indicate that the explainer is more truthful to the rule $c$. Observe that correctness is essentially the precision with which an explainer returns rule followers and breakers. Finally, we note that an important assumption is that the model $f$ is itself truthful to the rule, i.e., it has correctly learned the rule $c$, a condition we can check after training.
@@ 4.1 Relevance
We employ cosine similarity to compute relevance, given that the dataset comprises image embeddings produced by a pre-trained model. Utilizing the graph representation, we determine the average similarity between a explanandum node (explanandum) and its corresponding neighbors (example). The first column of Table 1, shows the explainer relevance. RIF demonstrates superior performance, whereas the remaining methods exhibit similar scores.
@@ 4.3 Correctness
Before evaluating the correctness of the explanation, it is crucial to ensure that the model accurately reflects the rule and has effectively learned it. Three metrics are used for this assessment: 1. Accuracy of Learning the Rule: This metric evaluates the model’s performance on test samples that correspond to the rule. 2. Log-Likelihood: It is expected that there will be a significant change in the log-likelihood of intervened points ($LL_{i}$) after the introduction of the rule, while the log-likelihood of untouched points ($LL_{u}$) should remain relatively stable. 3. Probability Scores: A notable change is anticipated in the probability scores of intervened points ($Ps_{i}$) compared to the untouched points ($Ps_{u}$). Table 2 illustrates the results of these metrics. In all cases, the model has successfully learned the rule without affecting its decisions for the untouched points.
Table 2: Model’s assessment in learning the rules
Acc $LL_{i}$ $LL_{u}$ $Ps_{i}$ $Ps_{u}$
Rule 1 0.83 Before -5.87 -9.4 100 15
After -0.42 -9.2
Rule 2 0.85 Before -12 -9.3 100 24.5
After -3.4 -7.2
Rule 3 0.92 Before -0.07 -10.6 98 12
After -1.83 -9.5
While explaining an explanandum following the rule we expect rule-followers in the training set with the same label to have a positive influence on the prediction, and rule-breakers with the opposite label to have a negative influence. Table 1 shows that IF and DataModels perform well in terms of correctness, but are inconsistent in relevance and distinguishability (domain and overlap), which is expected due to the presence of globally important class outliers. In contrast, RIF performs poorly in uncovering followers and breakers, because of its loss-based outlier elimination. RIF treats training data with high loss as undesired, and excludes them from explanation lists—the rationale is that such data would appear in all explanations and thus have little utility. But in this case, it is precisely the rule followers and particularly the minority of rule breakers that have high losses due to the ambiguity in the labeling rule. On the contrary, as previously discussed, the examples with very low loss are rewarded and appear in the explanations. TraceIn also fails to uncover the rule due to its low efficiency in identifying truly important examples, as demonstrated in [10].
@@ 5 Conclusion
Our analysis suggests that current example-based explainability techniques are susceptible to the influence of class outliers. Outright removal of outliers from explanations or datasets may not always be advisable, as they can possess valuable insights and provide explanations for similar examples encountered by models. Future research could focus on devising more robust explainability methods capable of extracting values from class outliers without compromising the quality of explanations using our proposed metrics.